# Simultaneous Motion And Noise Estimation with Event Cameras

### v1

@@ 4.2 Denoising Results
Denoising Performance on E-MLB. Table 1 shows the benchmark results on the E-MLB dataset using the MESR metric. We categorize prior work into model-based methods and learning-based methods for convenience. Compared with other model-based techniques, our method ranks first or second, showcasing the efficacy of the proposed method on denoising. Even compared with learning-based ones, our method achieves competitive scores (e.g., second in E-MLB (Day) ND4–ND64 conditions, etc). Notice that, however, learning-based approaches utilize additional information (GT labels) during training, while our method does not. Nonetheless, the results demonstrate the efficacy of our approach, which contributes to the theoretical foundation for the denoising methodology of event data. We discuss further details of the evaluation metrics (ESR) in Sec. 7.
@@
dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ $5$ Hz Noise CMax [17] 20.000 709 6820.0007096820.000\,709\,6820.000 709 68 1.258 796 0971.2587960971.258\,796\,0971.258 796 097 124.641 377124.641377124.641\,377124.641 377 1.128 529 5931.1285295931.128\,529\,5931.128 529 593 – w/ Init. 8.274 876 2458.2748762458.274\,876\,2458.274 876 245 1.273 697 0451.2736970451.273\,697\,0451.273 697 045 20.659 240 3320.6592403320.659\,240\,3320.659 240 33 1.214 159 4591.2141594591.214\,159\,4591.214 159 459 Downsampling (Best) 8.807 703 958.807703958.807\,703\,958.807 703 95 1.272 907 8221.2729078221.272\,907\,8221.272 907 822 124.868 639 7124.8686397124.868\,639\,7124.868 639 7 1.107 758 3951.1077583951.107\,758\,3951.107 758 395 – w/ Init. 8.226 219 4448.2262194448.226\,219\,4448.226 219 444 1.273 725 3071.2737253071.273\,725\,3071.273 725 307 20.618 615 8620.6186158620.618\,615\,8620.618 615 86 1.214 188 4551.2141884551.214\,188\,4551.214 188 455 Ours (Best) 8.521 574 0738.5215740738.521\,574\,0738.521 574 073 1.273 446 2961.2734462961.273\,446\,2961.273 446 296 97.584 574 8497.5845748497.584\,574\,8497.584 574 84 1.140 880 661.140880661.140\,880\,661.140 880 66 – w/ Init. $8.170$ 1.273 725 2991.2737252991.273\,725\,2991.273 725 299 20.604 334 6220.6043346220.604\,334\,6220.604 334 62 1.214 179 0181.2141790181.214\,179\,0181.214 179 018 BAF 19.674 893 0519.6748930519.674\,893\,0519.674 893 05 1.259 772 491.259772491.259\,772\,491.259 772 49 125.028 145 1125.0281451125.028\,145\,1125.028 145 1 1.127 308 9021.1273089021.127\,308\,9021.127 308 902 – w/ Init. 8.253 078 2028.2530782028.253\,078\,2028.253 078 202 1.273 636 991.273636991.273\,636\,991.273 636 99 20.550 436 620.550436620.550\,436\,620.550 436 6 1.213 505 481.213505481.213\,505\,481.213 505 48
@@ 4.3 Joint Motion Estimation Results
Ego-motion on ECD. One of the merits of the proposed joint estimation method is that it can improve motion estimation results. Here, the results of ego-motion (e.g., angular velocity) estimation are displayed in Fig. 5. We used the classical ECD dataset and injected BA noise from the DND21 recordings with different noise rates. As no prior work simultaneously estimates motion and noise, we compare against sequential filters: random sampling (of the input data with a uniform distribution) and BAF [9]. As shown in the IWEs, our method produces the sharpest edges (cf. (b) and (c)) and the most reasonable signal-noise classification among others (cf. (a), (b), and (c)). Note that the original ECD recordings have noise events besides the ones that are injected using DND21 recordings, which makes the calculation of ROC (similar to Sec. 3.2) intricate. Nonetheless, we report ROC results in the supplementary.
@@
Optical flow on MVSEC. The results on more complex motion estimation (e.g., optical flow) are shown in Tab. 4, using the MVSEC dataset [61], with BA noise from the DND21 dataset. Due to the $5$Hz BA noise, the baseline method MultiCM [47] degrades. Our denoising method improves the flow accuracy, showcasing the efficacy of simultaneous noise estimation. A challenge of evaluating denoising in these real-world datasets (e.g., ECD, MVSEC) is the existing noise in the original sequences. Hence, we only evaluate motion estimation metrics, not ROC or AUC.
indoor1 indoor2 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [47] 2.675 668 1172.6756681172.675\,668\,1172.675 668 117 32.098 464 8132.0984648132.098\,464\,8132.098 464 81 2.746 023 6482.7460236482.746\,023\,6482.746 023 648 33.474 068 5733.4740685733.474\,068\,5733.474 068 57 Ours 2.516 871 5322.5168715322.516\,871\,5322.516 871 532 29.475 340 6529.4753406529.475\,340\,6529.475 340 65 2.695 166 2252.6951662252.695\,166\,2252.695 166 225 33.046 739 9333.0467399333.046\,739\,9333.046 739 93 indoor3 outdoor1 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [47] 2.731 729 672.731729672.731\,729\,672.731 729 67 33.218 937 4433.2189374433.218\,937\,4433.218 937 44 1.960 064 5411.9600645411.960\,064\,5411.960 064 541 18.159 503 3618.1595033618.159\,503\,3618.159 503 36 Ours 2.653 875 2772.6538752772.653\,875\,2772.653 875 277 32.241 488 8832.2414888832.241\,488\,8832.241 488 88 1.864 629 221.864629221.864\,629\,221.864 629 22 17.554 100 1417.5541001417.554\,100\,1417.554 100 14
Table 4: Optical flow estimation results on MVSEC [61] ($dt=4$ frames, i.e., 89 ms) with BA noise at $5$ Hz.
@@ 4.6 Ablation of Motion Estimator
The proposed approach can be combined with other estimation methods, such as deep-neural–network (DNN) flow estimators. To this end, we show an example of E-RAFT [21]. The model-based method alone [47] is known to output a degraded flow in night sequences with lots of noise, such as zurich_city_12a [20]. As shown in Fig. 10, our method can identify and discard noisy events. E-RAFT provides sharper and more natural edges in IWEs after denoising compared to MultiCM [47] (thanks to its more accurate flow). ESR scores are similar: $2.39$ for E-RAFT and $2.45$ for MultiCM. While we focus on the inference of DNN in the proposed pipeline, future research could look into accommodating training within the proposed joint estimation framework.
@@ 4.7 Runtime
Using $30$k events from the ECD dataset ($240\times 180$ px) on a Mac M1 CPU (2020), the denoising operations (scoring and ranking) take less than $0.1$ s extra computation in total, out of 2 s of the entire processing time (i.e., $5$% computation increase with respect to the base method). Nonetheless, accommodating learning-based approaches in our framework could be key to the trade-off between accuracy and speed.
@@
(a) Signal Events (b) IWE (b) Flow Estimation
Figure 10: Effect of adopting two different motion estimators (E-RAFT or MultiCM) on the DSEC night scene (zurich_12a) [20].
@@ 5 Conclusion
Motion is an intrinsic part of event camera data, since scene edges cannot be sensed without it. We propose the first joint estimation framework for denoising and generic motion estimation, which connects two topics that have traditionally been disconnected in event-based vision. The experiments show that the proposed method achieves state-of-the-art results on the E-MLB benchmark and competitive results on the DND21 benchmark for denoising, while showing its efficacy on motion estimation and intensity reconstruction tasks. Moreover, the proposed method is flexible, as it allows replacing the 1-step motion estimation of CMax with any other motion estimator, such as DNNs. We believe that the proposed approach contributes not only to strengthening the theory of event-data denoising but also to expanding practical denoising use-cases via the code published.
@@ Acknowledgments
We would like to thank Dr. B. Jiang for useful discussions. This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
@@ 6.1 Video
We encourage readers to inspect the attached video, which summarizes the method and the results.
$\text{iter}=0$
@@
Figure 11: Evolutions of signal, noise, and motion (IWE) during optimization. The edge structure (e.g., green boxes in (b) Noise) converges to move to signal events, while the CMax converges to the sharp IWE (i.e., expected motion parameters).
@@
1Hz 3Hz 5Hz 7Hz 10Hz hotel driving hotel driving hotel driving hotel driving hotel driving Model-based KNoise [30] $0.6773$ $0.6296$ $0.6521$ $0.6230$ $0.6703$ $0.6235$ $0.6583$ $0.6164$ $0.6413$ $0.6142$ DWF [24] $0.9268$ $0.7409$ $0.8930$ $0.7099$ $0.8620$ $0.6901$ $0.8338$ $0.6747$ $0.7958$ $0.6563$ BAF [9] $0.9535$ $0.8479$ $0.9197$ $0.8155$ $0.8916$ $0.7930$ $0.8662$ $0.7732$ $0.8366$ $0.7479$ Ynoise [14] $0.9690$ $0.9409$ $0.9517$ $0.9240$ $0.9234$ $0.9093$ $0.9177$ $0.8972$ $0.8987$ $0.8800$ TS [33] $0.9716$ $0.9307$ $0.9721$ $0.9260$ $0.9606$ $0.9270$ $0.9654$ $0.9241$ $0.9620$ $0.9202$ Ours 1.013 802 194 317 176 21.01380219431717621.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.88211969313369480.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.967 937 633 567 141 90.96793763356714190.967\,937\,633\,567\,141\,90.967 937 633 567 141 9 0.850 838 015 623 797 50.85083801562379750.850\,838\,015\,623\,797\,50.850 838 015 623 797 5 0.962 515 815 619 444 40.96251581561944440.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.85539848463811620.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.951 406 776 235 535 30.95140677623553530.951\,406\,776\,235\,535\,30.951 406 776 235 535 3 0.846 774 906 545 3740.8467749065453740.846\,774\,906\,545\,3740.846 774 906 545 374 0.960 564 724 999 644 20.96056472499964420.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.8362735551022360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] $0.9573$ $0.8873$ $0.9371$ $0.8771$ $0.9365$ $0.8748$ $0.9254$ $0.8654$ $0.9006$ $0.874$ MLPF [24] $0.9704$ $0.8887$ $0.9718$ $0.8873$ $0.9704$ $0.8845$ $0.9691$ $0.8817$ $0.9634$ $0.8761$ EDformer [28] $0.9928$ $0.9541$ $0.9891$ $0.9472$ $0.9845$ $0.9424$ $0.9792$ $0.9343$ $0.9699$ $0.9264$

### v2

@@ 4.2 Denoising Results
Denoising Performance on E-MLB. Table 1 shows the benchmark results on the E-MLB dataset using the MESR metric. We categorize prior work into model-based methods and learning-based methods for convenience. Compared with other model-based techniques, our method ranks first or second, showcasing the efficacy of the proposed method on denoising. Even compared with learning-based ones, our method achieves competitive scores (e.g., second in E-MLB (Day) ND4–ND64 conditions, etc). Notice that, however, learning-based approaches utilize additional information (GT labels) during training, while our method does not. Nonetheless, the results demonstrate the efficacy of our approach, which contributes to the theoretical foundation of event-denoising methodologies. We discuss further details of the evaluation metrics (ESR) in Sec. 5.
@@
dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ 555 Hz Noise CMax [17] 20.000 709 6820.000\,709\,6820.000 709 68 1.258 796 0971.258\,796\,0971.258 796 097 124.641 377124.641\,377124.641 377 1.128 529 5931.128\,529\,5931.128 529 593 – w/ Init. 8.274 876 2458.274\,876\,2458.274 876 245 1.273 697 0451.273\,697\,0451.273 697 045 20.659 240 3320.659\,240\,3320.659 240 33 1.214 159 4591.214\,159\,4591.214 159 459 Downsampling (Best) 8.807 703 958.807\,703\,958.807 703 95 1.272 907 8221.272\,907\,8221.272 907 822 124.868 639 7124.868\,639\,7124.868 639 7 1.107 758 3951.107\,758\,3951.107 758 395 – w/ Init. 8.226 219 4448.226\,219\,4448.226 219 444 1.273 725 3071.273\,725\,3071.273 725 307 20.618 615 8620.618\,615\,8620.618 615 86 1.214 188 4551.214\,188\,4551.214 188 455 Ours (Best) 8.521 574 0738.521\,574\,0738.521 574 073 1.273 446 2961.273\,446\,2961.273 446 296 97.584 574 8497.584\,574\,8497.584 574 84 1.140 880 661.140\,880\,661.140 880 66 – w/ Init. 8.1708.1708.170 1.273 725 2991.273\,725\,2991.273 725 299 20.604 334 6220.604\,334\,6220.604 334 62 1.214 179 0181.214\,179\,0181.214 179 018 BAF 19.674 893 0519.674\,893\,0519.674 893 05 1.259 772 491.259\,772\,491.259 772 49 125.028 145 1125.028\,145\,1125.028 145 1 1.127 308 9021.127\,308\,9021.127 308 902 – w/ Init. 8.253 078 2028.253\,078\,2028.253 078 202 1.273 636 991.273\,636\,991.273 636 99 20.550 436 620.550\,436\,620.550 436 6 1.213 505 481.213\,505\,481.213 505 48
@@ 4.3 Joint Motion Estimation Results
Ego-Motion on ECD. One of the merits of the proposed joint estimation method is that it can improve motion estimation results. Here, the results of ego-motion (e.g., angular velocity) estimation are displayed in Fig. 5. We used the classical ECD dataset and injected BA noise from the DND21 recordings with different noise rates. As no prior work simultaneously estimates motion and noise, we compare against sequential filters: random sampling (of the input data with a uniform distribution) and BAF [9] that is the most widely-used denoising method to date. As shown in the IWEs, our method produces the sharpest edges (cf. (b) and (c)) and the most reasonable signal-noise classification among others (cf. (a), (b), and (c)). Note that the original ECD recordings have noise events besides the ones that are injected using DND21 recordings, which makes the calculation of ROC (similar to Sec. 3.2) intricate. Nonetheless, we report ROC results in the supplementary.
@@
Optical Flow on MVSEC. The results on more complex motion estimation (e.g., optical flow) are shown in Tab. 4, using the MVSEC dataset [63], with BA noise from the DND21 dataset. Due to the  [^5] Hz BA noise, the baseline method MultiCM [48] degrades. Our denoising method improves the flow accuracy, showcasing the efficacy of simultaneous noise estimation. A challenge of evaluating denoising in these real-world datasets (e.g., ECD, MVSEC) is the existing noise in the original sequences. Hence, we only evaluate motion estimation metrics, not ROC or AUC.
indoor1 indoor2 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [48] 2.675 668 1172.675\,668\,1172.675 668 117 32.098 464 8132.098\,464\,8132.098 464 81 2.746 023 6482.746\,023\,6482.746 023 648 33.474 068 5733.474\,068\,5733.474 068 57 Ours 2.516 871 5322.516\,871\,5322.516 871 532 29.475 340 6529.475\,340\,6529.475 340 65 2.695 166 2252.695\,166\,2252.695 166 225 33.046 739 9333.046\,739\,9333.046 739 93 indoor3 outdoor1 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [48] 2.731 729 672.731\,729\,672.731 729 67 33.218 937 4433.218\,937\,4433.218 937 44 1.960 064 5411.960\,064\,5411.960 064 541 18.159 503 3618.159\,503\,3618.159 503 36 Ours 2.653 875 2772.653\,875\,2772.653 875 277 32.241 488 8832.241\,488\,8832.241 488 88 1.864 629 221.864\,629\,221.864 629 22 17.554 100 1417.554\,100\,1417.554 100 14
Table 4: Optical flow estimation results on MVSEC [63] ($dt=4$ frames, i.e., 89 ms) with BA noise at 555 Hz.
@@ 4.6 Ablation of Motion Estimator
The proposed approach can be combined with other estimation methods, such as deep-neural–network (DNN) flow estimators. To this end, we show an example of E-RAFT [21]. The model-based method alone [48] is known to output a degraded flow in night sequences with lots of noise, such as zurich_city_12a [20]. As shown in Fig. 10, our method can identify and discard noisy events. E-RAFT provides sharper and more natural edges in IWEs after denoising compared to MultiCM [48] (thanks to its more accurate flow). ESR scores are similar: 2.392.392.39 for E-RAFT and 2.452.452.45 for MultiCM. While we focus on the inference of DNN in the proposed pipeline, future research could look into accommodating training within the proposed joint estimation framework.
@@ 4.7 Runtime
Using 303030k events from the ECD dataset (240×180240$\times$ 180240 × 180 px) on a Mac M1 CPU (2020), the denoising operations (scoring and ranking) take less than 0.10.10.1 s extra computation in total, out of 2 s of the entire processing time (i.e., 555% computation increase with respect to the base method). Nonetheless, accommodating learning-based approaches in our framework could be key to the trade-off between accuracy and speed.
@@
(a) Signal Events (b) IWE (c) Flow Estimation
Figure 10: Effect of adopting two different motion estimators (E-RAFT or MultiCM) on the DSEC night scene (zurich_12a) [20]. Despite small IWE differences (e.g., in the car’s hood), the proposed framework supports several motion estimators.
5 Limitations
Our method expects as parameter a target noise level, but the intrinsic (“true”) noise level depends on the camera, scene, etc. and is therefore unknown. More research in this direction is needed for automatic parameter tuning. Also, notice that there might be a potential limitation of the denoising metric ESR, since it depends on the motion parameter (see Suppl. Mat. of [29]). Although our denoising efficacy is confirmed from multiple perspectives, we believe that advancing denoising metrics for real-world data (where it gets interesting due to the absence of GT labels) is a central direction for future work in event-based vision.
@@ 6 Conclusion
In summary, this work proposes that: ($i$) Joint estimation of noise and motion is a novel problem setting compared to conventional (isolated) denoising, and is sensitive given the event data’s inherent motion dependency. ($ii$) The proposed joint estimation framework can accommodate various motion estimators and has shown its efficacy in improving both denoising and motion estimation accuracies. ($iii$) It is important to evaluate denoising results on both, denoising benchmarks and downstream tasks (e.g., motion estimation, intensity reconstruction), since current denoising benchmarks have shortcomings in terms of metrics, data, and baselines. In particular, our method achieves notable denoising results (state-of-the-art on the E-MLB benchmark and competitive on DND21) while demonstrating effectiveness across motion estimation and intensity reconstruction tasks. Our approach advances both event-denoising theory and practical denoising use-cases.
@@ Acknowledgments
We would like to thank Dr. B. Jiang for useful discussions. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
@@ 7.1 Video
We encourage readers to inspect the attached video, which summarizes the method and results.
[…]
@@
Figure 11: Evolutions of signal, noise, and motion (IWE) during optimization. The edge structure (e.g., green boxes in (b) Noise) converges to move to signal events, while CMax converges to the sharp IWE (i.e., expected motion parameters).
@@
1Hz 3Hz 5Hz 7Hz 10Hz hotel driving hotel driving hotel driving hotel driving hotel driving Model-based BAF [9] 0.95350.95350.9535 0.84790.84790.8479 0.91970.91970.9197 0.81550.81550.8155 0.89160.89160.8916 0.79300.79300.7930 0.86620.86620.8662 0.77320.77320.7732 0.83660.83660.8366 0.74790.74790.7479 TS [34] 0.97160.97160.9716 0.93070.93070.9307 0.97210.97210.9721 0.92600.92600.9260 0.96060.96060.9606 0.92700.92700.9270 0.96540.96540.9654 0.92410.92410.9241 0.96200.96200.9620 0.92020.92020.9202 KNoise [31] 0.67730.67730.6773 0.62960.62960.6296 0.65210.65210.6521 0.62300.62300.6230 0.67030.67030.6703 0.62350.62350.6235 0.65830.65830.6583 0.61640.61640.6164 0.64130.64130.6413 0.61420.61420.6142 Ynoise [14] 0.96900.96900.9690 0.94090.94090.9409 0.95170.95170.9517 0.92400.92400.9240 0.92340.92340.9234 0.90930.90930.9093 0.91770.91770.9177 0.89720.89720.8972 0.89870.89870.8987 0.88000.88000.8800 DWF [24] 0.92680.92680.9268 0.74090.74090.7409 0.89300.89300.8930 0.70990.70990.7099 0.86200.86200.8620 0.69010.69010.6901 0.83380.83380.8338 0.67470.67470.6747 0.79580.79580.7958 0.65630.65630.6563 Ours 1.013 802 194 317 176 21.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.967 937 633 567 141 90.967\,937\,633\,567\,141\,90.967 937 633 567 141 9 0.850 838 015 623 797 50.850\,838\,015\,623\,797\,50.850 838 015 623 797 5 0.962 515 815 619 444 40.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.951 406 776 235 535 30.951\,406\,776\,235\,535\,30.951 406 776 235 535 3 0.846 774 906 545 3740.846\,774\,906\,545\,3740.846 774 906 545 374 0.960 564 724 999 644 20.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] 0.95730.95730.9573 0.88730.88730.8873 0.93710.93710.9371 0.87710.87710.8771 0.93650.93650.9365 0.87480.87480.8748 0.92540.92540.9254 0.86540.86540.8654 0.90060.90060.9006 0.8740.8740.874 MLPF [24] 0.97040.97040.9704 0.88870.88870.8887 0.97180.97180.9718 0.88730.88730.8873 0.97040.97040.9704 0.88450.88450.8845 0.96910.96910.9691 0.88170.88170.8817 0.96340.96340.9634 0.87610.87610.8761 EDformer [29] 0.99280.99280.9928 0.95410.95410.9541 0.98910.98910.9891 0.94720.94720.9472 0.98450.98450.9845 0.94240.94240.9424 0.97920.97920.9792 0.93430.93430.9343 0.96990.96990.9699 0.92640.92640.9264